# for Large Language Models

### v1

@@ 6 Conclusion
In this work, we introduce BenchMAX, a comprehensive, high-quality, and multiway parallel multilingual benchmark comprising ten tasks designed to assess crucial capabilities across seventeen diverse languages. The multilingual task data is initially translated from English using machine translation and subsequently refined through multiple iterations of post-editing by native speakers, ensuring high data quality. Through extensive model evaluations, we find that the language-agnostic capabilities of state-of-the-art LLMs remain uneven across different languages. While increasing model size consistently enhances multilingual performance, the performance gap between English and other languages persists, highlighting the need for further efforts to achieve balanced multilingual capabilities.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Table 8: Ranks of four models in our tasks compared to discriminative tasks. The rankings of Discriminative tasks are from Dang et al. (2024).
Model Rule-based Func Compl. Discriminative
Qwen2.5-7B No.1 No.1 No.4
Llama3.1-8B No.3 No.3 No.3
Aya-Expanse-8B No.4 No.4 No.2
Gemma2-9B No.2 No.2 No.1
@@ Appendix A Capability and Task Data Selection
  • Rule-based Intruction Following: We collect data from IFEval (Zhou et al., 2023), which is a benchmark for evaluating the instruction following abilities of LLMs, composed of around 500 verifiable instructions and can be evaluated for accuracy using automated rules. Note that the accuracy for IFEval is the average of the four accuracies (i.e. prompt-strict, prompt-loose, inst-strict and inst-loose accuracies), following (Dubey et al., 2024).
  • Model-based Instruction Following: We collect data from Arena-hard (Li et al., 2024) which contains 500 real-world instructions from the Chatbot Arena (Chiang et al., 2024), and m-ArenaHard [^6] https://huggingface.co/datasets/CohereForAI/m-ArenaHard which contains translated multilingual versions. This benchmark can provide better model separability and higher alignment with human preference. It is assessed by the Win Rate of the testing model in comparison to the baseline model, GPT-4o, judged against GPT-4o-mini.
@@
  • Function Completion: We collect data from Humaneval+ (Liu et al., 2024) which is an augmented version of HumanEval (Chen et al., 2021), comprising an expanded test cases. Each problem in the benchmark gives a definition of a Python function accompanied by an English docstring, and requires LLMs to complete the function.
  • Problem Solving: We collect data from LiveCodeBench  [^7] We adopt the code generation subset in LiveCodeBench v4 as the original English dataset. (Jain et al., 2024) which provides a more rigorous assessment of the code generation capabilities. It is a much harder benchmark by collecting coding problems in natural language from real competition platforms with live updates.
@@
involves understanding and generating coherent text from extensive input sequences, allowing the model to capture dependencies and relationships within lengthy texts. This paper focuses on the long-context evaluation of multilingual settings based on the RULER benchmark (Hsieh et al., 2024).
@@
encompasses thinking logically, drawing conclusions, making inferences, and solving problems by processing data, applying rules, and utilizing various forms of logic and knowledge representation. Pushing LLMs beyond surface-level tasks, we extend MGSM (Shi et al., 2023) and GPQA (Rein et al., 2023) requiring deeper understanding and reasoning across different context.
@@
requires the model to translate user queries into executable functions for calling in operating software tools. We extend Nexus (Srinivasan et al., 2023) to a multilingual version, which is adopted by Llama3 (Dubey et al., 2024).
@@
  • General: General domain data are composed of Flores-200 (Costa-jussà et al., 2022), TED (Cettolo et al., 2012) and WMT24 (Kocmi et al., 2024) testsets. In BenchMAX, we include parallel data from 17 selected languages.
@@ B.7 Long-Context Question Answering Dataset
The haystacks, needles, paragraphs and questions related to QAs are translated to other languages. We use the parallel testsets from the UN corpus (Ziemski et al., 2016) as the haystack. The English version contains about 128k tokens, and we extend it to other languages using Google Translate. The sentence of the needle is also translated into 16 other languages, in which UUIDs are employed as keys and values that are not translated. With respect to the QA data, we translate the paragraphs and questions in XQuAD (Artetxe et al., 2020) to the languages we need. Note that we also use the trick in translating IFEval to extract the answer spans. With access to our multilingual haystacks, needles and paragraphs, we are able to synthesize the multilingual long-context testsets.
@@ Appendix C Model Information
Llama3.1-Instruct (Dubey et al., 2024)
@@
Qwen2.5-Instruct (Qwen Team, 2024)
is a collection of multilingual language models with several sizes, ranging from 0.5B to 72B. The models are trained with multilingual tokens in both pre-training stage and post-training stage, and are rigorously evaluated on several multilingual tasks. In our experiments, we evaluate the 7B version and the 72B version of Qwen2.5-Instruct.
Aya-Expanse (Dang et al., 2024)
@@
InternLM2.5-chat (Cai et al., 2024)
is the successor of InternLM (Team, 2023), which is claimed as a multilingual model. We include the 7B version and 20B version in our experiments. InternLM2.5-7B-chat-1m is a long-context variant supporting context windows with 1M tokens.
DeepSeek-V3 (DeepSeek-AI et al., 2024)
@@
GPT-4o & GPT-4o-mini (OpenAI, 2024)
@@ Appendix D Details about Prompt Templates
Table 13: The prompt templates of the long-context modelling task.
@@
Table 14: The GEMBA-SQM prompt.
@@
Table 15: LLM-Judge Instruction
@@ 5. Human B is significantly better: [[B>>much-greater-than$>$ >A]]
Table 16: Prompt for translating the Function Completion task.
@@
Table 17: Prompt for translating the Problem Solving task.

### v2

@@ 6 Conclusion
In this work, we introduce BenchMAX, a comprehensive, high-quality, and parallel multilingual benchmark comprising ten tasks designed to assess crucial capabilities across seventeen diverse languages. The multilingual task data is initially translated from English using machine translation and subsequently refined through multiple iterations of post-editing by native speakers, ensuring high data quality. Through extensive model evaluations, we find that the language-agnostic capabilities of state-of-the-art LLMs remain uneven across different languages. While increasing model size consistently enhances multilingual performance, the performance gap between English and other languages persists, highlighting the need for further efforts to achieve balanced multilingual capabilities.
@@ Appendix A Capability and Task Data Selection
  • Rule-based Intruction Following: We collect data from IFEval Zhou et al. (2023), which is a benchmark for evaluating the instruction following abilities of LLMs, composed of around 500 verifiable instructions and can be evaluated for accuracy using automated rules. Note that the accuracy for IFEval is the average of the four accuracies (i.e. prompt-strict, prompt-loose, inst-strict and inst-loose accuracies), following Dubey et al. (2024).
  • Model-based Instruction Following: We collect data from Arena-hard Li et al. (2024) which contains 500 real-world instructions from the Chatbot Arena Chiang et al. (2024), and m-ArenaHard [^6] https://huggingface.co/datasets/CohereForAI/m-ArenaHard which contains translated multilingual versions. This benchmark can provide better model separability and higher alignment with human preference. It is assessed by the Win Rate of the testing model in comparison to the baseline model, GPT-4o, judged against DeepSeek-V3.
@@
  • Function Completion: We collect data from Humaneval+ Liu et al. (2024) which is an augmented version of HumanEval Chen et al. (2021), comprising an expanded test cases. Each problem in the benchmark gives a definition of a Python function accompanied by an English docstring, and requires LLMs to complete the function.
  • Problem Solving: We collect data from LiveCodeBench  [^7] We adopt the code generation subset in LiveCodeBench v4 as the original English dataset. Jain et al. (2024) which provides a more rigorous assessment of the code generation capabilities. It is a much harder benchmark by collecting coding problems in natural language from real competition platforms with live updates.
@@
involves understanding and generating coherent text from extensive input sequences, allowing the model to capture dependencies and relationships within lengthy texts. This paper focuses on the long-context evaluation of multilingual settings based on the RULER benchmark Hsieh et al. (2024).
@@
encompasses thinking logically, drawing conclusions, making inferences, and solving problems by processing data, applying rules, and utilizing various forms of logic and knowledge representation. Pushing LLMs beyond surface-level tasks, we extend MGSM Shi et al. (2023) and GPQA Rein et al. (2023) requiring deeper understanding and reasoning across different context.
@@
requires the model to translate user queries into executable functions for calling in operating software tools. We extend Nexus Srinivasan et al. (2023) to a multilingual version, which is adopted by Llama3 Dubey et al. (2024).
@@
  • General: General domain data are composed of Flores-200 Costa-jussà et al. (2022), TED Cettolo et al. (2012) and WMT24 Kocmi et al. (2024) testsets. In BenchMAX, we include parallel data from 17 selected languages.
@@ B.7 Long-Context Question Answering Dataset
The haystacks, needles, paragraphs and questions related to QAs are translated to other languages. We use the parallel testsets from the UN corpus Ziemski et al. (2016) as the haystack. The English version contains about 128k tokens, and we extend it to other languages using Google Translate. The sentence of the needle is also translated into 16 other languages, in which UUIDs are employed as keys and values that are not translated. With respect to the QA data, we translate the paragraphs and questions in XQuAD Artetxe et al. (2020) to the languages we need. Note that we also use the trick in translating IFEval to extract the answer spans. With access to our multilingual haystacks, needles and paragraphs, we are able to synthesize the multilingual long-context testsets.
@@ Appendix C Model Information
Llama3.1-Instruct Dubey et al. (2024)
@@
Qwen2.5-Instruct Qwen Team (2024)
is a collection of multilingual language models with several sizes, ranging from 0.5B to 72B. The models are trained with multilingual tokens in both pre-training stage and post-training stage, and are rigorously evaluated on several multilingual tasks. In our experiments, we evaluate the 7B, 32B ,and 72B version of Qwen2.5-Instruct.
Aya-Expanse Dang et al. (2024)
@@
InternLM2.5-chat Cai et al. (2024)
is the successor of InternLM Team (2023), which is claimed as a multilingual model. We include the 7B version and 20B version in our experiments. InternLM2.5-7B-chat-1m is a long-context variant supporting context windows with 1M tokens.
DeepSeek-V3 DeepSeek-AI et al. (2024)
@@
DeepSeek-R1-Distill-Llama & DeepSeek-R1-Distill-Qwen
Guo et al. (2025) are dense models with long reasoning capabilities, and are distilled from DeepSeek-R1 based on Llama3.1-8B, Llama3.3-70B-Instruct, Qwen2.5-Math-7B, and Qwen2.5-32B.
GPT-4o & GPT-4o-mini OpenAI (2024)
@@ Appendix D Details about Prompt Templates
Table 15: The prompt templates of the long-context modelling task.
@@
Table 16: The GEMBA-SQM prompt.
@@ 5. Human B is significantly better: [[B>>much-greater-than$>$ >A]]
Table 17: LLM-Judge Instruction
@@
Table 18: Prompt for translating the Function Completion task.
@@
Table 19: Prompt for translating the Problem Solving task.